Ericka is twice as old as her sister and half as old as her mother. And in 22 years, her sister will be half as old as the mother. How old is Ericka now?

Answer by ewatrrr(24785)


Hi
Ericka is twice as old as her sister and half as old as her mother
Let x, x/2 and x/4 represent the mother's, ericka's and her sister ages.
Write as we read***
(x/4 + 22) = 1/2(x+22)
Solving for x
x/4 + 22 = 1/2x + 11
11 = x/4
x = 44, the mother's age now, ericka is 22 and her sister 11
CHECKING our Anwer***
11 + 22 = 33 = (44+22)/2